# Waffle: Multi-Modal Model for Automated Front-End Development

### v1

@@ 4.3 Structure-Aware Attention’s Effect
To demonstrate how structure-aware attention helps MLLMs focus on the correct elements (like parent and sibling elements) during generation, we simulate 20 cases where generation errors occur mid-process. Ideally, mistakes within sibling elements should not disrupt the generation of the main element. We select 20 high-performing samples generated by VLM-WebSight fine-tuned with Waffle and Waffle-${}_{\texttt{-attn}}$. Using the HTML code generated by the models, we manually crop and mutate sections to simulate errors. The models are then tasked with re-completing the HTML code starting from the error. Our goal is to evaluate whether the models are robust enough to handle intermediate errors without causing major failures in subsequent generations.
@@
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline=false] Summary: Integrating structure-aware attention brings great stability to model generation, making models more robust against intermediate mistakes, and reducing the performance drop by 23.31%, from 27.55% to 4.24%.
@@ 5 Related Work
5.1 Multi-Model Large Language Models
Recent advances in vision and language models have greatly improved MLLMs’ capabilities in tasks like image captioning Radford et al. (2021); Zhai et al. (2023); Li et al. (2022, 2023b); McKinzie et al. (2024); Alayrac et al. (2022); Laurençon et al. (2023), text-to-image generation Ramesh et al. (2022); Rombach et al. (2021), and visual question answering Dai et al. (2023); Liu et al. (2023b, a); Bai et al. (2023). While popular models like Llava Liu et al. (2023b, a), Qwen-VL Bai et al. (2023), and Vary Wei et al. (2023) perform well in general image tasks, they don’t focus on converting UI images to HTML code. To address this, we propose Waffle, a fine-tuning method that equips MLLMs with the domain-specific knowledge needed for UI-to-HTML generation.
@@ 6 Limitation
One limitation of Waffle is that it has only been implemented on two models: VLM-WebSight and Moondream2. While Waffle could potentially be applied to any MLLM, further exploration is limited by computing resources. Our experiments show that Waffle brings significant improvements over standard fine-tuning on these two models, indicating some level of generalizability. Another limitation is that the metrics we use do not fully capture human evaluation. HTML-Match overlooks CSS styling, and metrics like CW-SSIM, CLIP, and LLEM are similarity-based, which can lead to unreliable scores. Evaluating HTML code automatically is challenging, so we include CLIP and LLEM, as used in previous work Si et al. (2024); Gui et al. (2024), along with CW-SSIM and HTML-Match to ensure fair evaluation.
7 Conclusion
This work presents Waffle, a fine-tuning pipeline for UI-to-HTML code generation. Waffle introduces a structure-aware attention mechanism to capture HTML structure and employs contrastive learning to align visual and textual understanding, aiding MLLMs in distinguishing subtle webpage differences. Waffle outperforms standard fine-tuning on two backbone MLLMs, with improvements of up to 9 pp in HTML Match, 0.0982 in CW-SSIM, 32.99 in CLIP, and 27.12 pp in LLEM. Ablation studies confirm that both key components of Waffle contribute to better cross-modality understanding and more robust code generation. Notably, Waffle is model-independent and can enhance any MLLMs for UI-to-HTML code generation.
@@ A.1 Mutation Rules
A.2 Case Study
Figure 5 shows the generation results for one instance from WebSight-Test. The generated webpage of GPT-4o has a CW-SSIM of 0.1353, significantly lower than that of 0.3760 from VLM-WebSight under standard fine-tuning. On the other hand, the webpage generated by VLM-WebSight fine-tuned by Waffle reaches an almost perfect CW-SSIM score, 0.9995. This example shows the effectiveness of using Waffle on the UI image to HTML code task.
Figure 5: Example test instance from WebSight-Test dataset, with the generated images by GPT-4o, Standard FT, and Waffle.
@@ A.2 Case Study
A.4 Contrastive Learning’s Effect
Standard FT 1.3395 0.1027
Waffle-${}_{\texttt{-attn}}$ 0.8447 0.6244
@@ A.4 Contrastive Learning’s Effect
A.5 Infrastructure

### v2

@@ 4.4 Structure-Aware Attention’s Effect
To demonstrate how structure-aware attention helps MLLMs focus on the correct structural elements (such as parent and sibling elements) during generation, we introduce controlled errors in the generation process. Specifically, we select 20 high-performing samples—those with the highest CW-SSIM scores—generated by VLM-WebSight fine-tuned with Waffle and Waffle-${}_{\texttt{-attn}}$. These samples indicate cases where the models originally performed well on UI-to-code generation. Our goal is to evaluate whether models fine-tuned with structure-aware attention exhibit greater robustness in handling intermediate errors compared to those without it.
For each selected sample, we manually modify the partially generated HTML code by editing the sibling elements of the correct structure, following the example in Figure 1. These modifications alter the rendered output, simulating realistic errors that could occur during generation. The models are then tasked with re-completing the HTML code while being provided the same UI image with the partially generated, incorrect HTML as input. Since the modifications affect sibling elements rather than the primary structure, an ideal model should recover from these errors without cascading failures in subsequent generations.
@@
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline=false] Summary: Integrating structure-aware attention brings stability to model generation, making models more robust against intermediate mistakes, and reducing the performance drop by 23.31%, from 27.55% to 4.24%, ensuring more consistent generation quality.
4.5 Contrastive Learning’s Effect
Standard FT 1.3395 0.1027
Waffle-${}_{\texttt{-attn}}$ 0.8447 0.6244
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline=false] Summary: Fine-tuning with contrastive learning (Waffle-${}_{\texttt{-attn}}$) enhances MLLM visual-textual alignment, making image and code embeddings more cohesive, reducing Euclidean distance while elevating the cosine similarity between them.
@@ 5 Related Work
5.1 Multi-Modal Large Language Models
Recent advances in vision and language models have greatly improved MLLMs’ capabilities in tasks like image captioning Zhai et al. (2023); Li et al. (2022, 2023b); Laurençon et al. (2023), text-to-image generation Ramesh et al. (2022); Rombach et al. (2021), visual question answering Liu et al. (2023b, a); Bai et al. (2023), and document editing Suri et al. (2024). While popular models like Llava Liu et al. (2023b, a), Qwen-VL Bai et al. (2023), and Vary Wei et al. (2023) perform well in general image tasks, they don’t focus on converting UI images to HTML code. To address this, we propose Waffle, a fine-tuning method that equips MLLMs with domain-specific knowledge needed for UI-to-HTML generation.
@@ 5.3 UI to HTML Generation
6 Conclusion
This work presents Waffle, a fine-tuning pipeline for UI-to-HTML code generation, that is generalizable to any transformer-based MLLMs. Waffle introduces a structure-aware attention mechanism to capture HTML structure and employs contrastive learning to align visual and textual understanding, aiding MLLMs in distinguishing subtle webpage differences. Waffle outperforms standard fine-tuning on two backbone MLLMs, with improvements of up to 9 pp in HTML Match, 0.0982 in CW-SSIM, 32.99 in CLIP, and 27.12 pp in LLEM. Ablation studies confirm that both key components of Waffle contribute to better cross-modality understanding and more robust code generation. Notably, Waffle is model-independent and can enhance any MLLMs for UI-to-HTML code generation.
@@ 7 Limitation
One limitation of Waffle is that it has only been implemented on two models: VLM-WebSight and Moondream2. While Waffle could potentially be applied to any MLLM, like Design2Code, further exploration is limited by computing resources. Our experiments show that Waffle brings significant improvements over standard fine-tuning on these two models, indicating some level of generalizability. Another limitation is that the metrics we use do not fully capture human evaluation. HTML-Match overlooks CSS styling, and metrics like CW-SSIM, CLIP, and LLEM are similarity-based, which can lead to unreliable scores. Evaluating HTML code automatically is challenging, so we include CLIP and LLEM, as used in previous work Si et al. (2024); Gui et al. (2024), along with CW-SSIM and HTML-Match to ensure fair evaluation.
8 Acknowledgements
We thank the anonymous reviewers for their feedback on this work. This research was supported in part by NSF 1901242 and 2006688 and a CFI fund. This work also used Anvil at Purdue University through allocation CIS240304 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program Boerner et al. (2023), which is supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
@@ A.1 Mutation Rules
A.2 Illustrating Example
@@
A.4 Additional Ablation of Contrastive Learning’s effect
@@
A.5 Attention Visualization
Figure 9: Selected Attention Weight Visualization of Waffle on a simplified example.
The visualization of attention weights in Figure 9 provides insight into the behavior of structure-aware attention and standard attention mechanisms. Four representative examples are selected from different attention heads, each demonstrating distinct characteristics.
The attention in (a) exhibits a sparse and discrete distribution while effectively leveraging HTML domain knowledge. Notably, elements such as <div id = "rightCol"> do not attend to the children of their siblings, allowing for a more efficient allocation of attention resources. This pattern suggests that structure-aware attention successfully encodes hierarchical relationships, prioritizing relevant dependencies.
The attention in (b) demonstrates a diagonalized pattern, aligning with the common attention sink phenomenon observed in deeper layers. The attention weights primarily concentrate along the diagonals, indicating that elements predominantly attend to themselves or closely positioned tokens. Despite this, the model retains structural awareness, as elements do not indiscriminately attend to siblings’ children. This behavior suggests that structure-aware attention maintains domain knowledge while refining local relationships in later layers.
[…]
The attention in (d) also exhibits a diagonalized pattern, indicative of attention sink behavior, where tokens predominantly attend to closely positioned elements. This pattern resembles the behavior observed in (b), yet lacks the structured constraints of structure-aware attention. The similarity between the two suggests that deeper layers across both models tend to favor local attention.
A.6 Waffle’s Computational Overhead
We discuss the computational overhead of Waffle in preprocessing, training, and inference separately.
Preprocessing The only additional one-time preprocessing cost incurred for each sample and tokenizer is the construction cost of the structure-aware attention masks. The construction process for these structure-aware attention masks exhibits a complexity of $O(B\cdot L)$, where $B$ denotes the buffer length in our specialized HTML parser used for tracking HTML tag tokens, and $L$ represents the overall token length. This compares to the $O(L)$ complexity for constructing a standard causal mask. Notably, the buffer $B$ solely contains tokens of currently incomplete HTML tags, rendering its size small relative to the full token length $L$. Furthermore, this preprocessing can be executed on a CPU and is exceptionally lightweight, as it does not involve matrix multiplication operations. Once these attention masks are generated, they are reused throughout the entire training process without incurring further overhead, making this initial computational investment negligible when amortized across the full training cycle.
Training The computational demands during the training phase of Waffle are influenced by two primary components: Structure-Aware Attention and Contrastive Learning.
The introduction of Structure-Aware Attention results in no additional computational cost during training. This is because the only modification pertains to the values within the attention mask, while the total number of operations remains the same to that of standard attention mechanisms.
Our implementation of Contrastive Learning, the image embeddings are computed once and subsequently reused for both standard fine-tuning tasks and the contrastive learning objectives. The main additional computation arises from generating text embeddings, which requires an extra forward pass through the decoder. Other minor computational elements, such as similarity calculations and the computation of the contrastive loss, are negligible in comparison to the overall training expenditure.
In our experimental configuration, utilizing 4x NVIDIA A100 GPUs with WebSight-VLM, standard fine-tuning necessitates approximately 26 hours. In contrast, training with Waffle requires about 34 hours, which constitutes a 31% increase. It is important to emphasize that this is a one-time training cost and does not adversely affect inference speed. We consider this increase a worthwhile investment, given the significant performance improvements achieved by Waffle.
Inference During inference, the additional computational cost introduced by Waffle stems from parsing the generated HTML and constructing the structure-aware attention mask for newly generated tokens. This construction process is identical to the one employed during preprocessing and, similarly, remains lightweight in comparison to the model’s other operational demands.
In conclusion, Waffle introduces a limited computational overhead during training and imposes minimal computation overhead during inference and preprocessing.
A.7 Analysis of Failure Examples
Figure 10: Figures of Design2Code task 42.
Figure 11: Figures of WebSight-Test task 45.
To provide a balanced evaluation of Waffle’s capabilities, this section presents an analysis of selected failure cases. We discuss two sets of illustrative examples, one from each test set (Design2Code and WebSight-Test), where the VLM-WebSight model fine-tuned by Waffle exhibited deviations from the ground truth.
Design2Code-42 The first case is from the Design2Code test set, example Design2Code-42, shown in Figure 10. The ground truth for this example is depicted in Figure 10 (a), while the output generated by our Waffle-enhanced VLM-WebSight model is shown in Figure 10 (b). For a comparative perspective, the result from VLM-WebSight under standard fine-tuning (SFT) is provided in Figure 10 (c).
Upon examining the output from Waffle, several discrepancies were identified when compared to the ground truth. A primary challenge appears to be the accurate recognition and reproduction of all textual elements. For instance, there was a noticeable difference in the domain name; the ground truth specifies "ESSAYSCRIBES.COM", whereas the generated version incorrectly rendered it as "EASYSCRIBES.COM". Further textual inaccuracies included differences in the primary heading or description on the first line, where "Essential Example" was expected but variations occurred, and section headings were altered, such as "How an Editorial Essay Works" in the ground truth becoming "How The Editor Writes an Essay" in the generated HTML.
Styling inconsistencies also emerged. The navigation menu in the Waffle-generated version exhibited different spacing and font treatment compared to the original design. Additionally, the navigation items themselves were simplified and featured slightly different labeling. Other minor errors were observed, such as in the footer area, where, although the copyright text was consistent, its formatting and color scheme diverged from the ground truth.
In summary, while VLM-WebSight fine-tuned by Waffle successfully captured the main structure of this real-world UI from the Design2Code test set, several minor errors in text rendering and styling persist. These indicate areas for further improvement in UI-to-HTML generation. It is pertinent to note, however, that these errors are considerably less severe than those observed in the output from the VLM-WebSight model under standard fine-tuning. The SFT version exhibited more drastic failures, with significant mistakes in color fidelity, overall structure, and text extraction.
WebSight-Test-45 The second failure case analyzed is WebSight-Test-45. The ground truth image for this example is available as Figure 11 (a), and the corresponding page generated by VLM-WebSight fine-tuned with Waffle is shown in Figure 11 (b).
In this instance, the generated output aligns well with the ground truth in terms of most visual elements. However, a key error was identified in the positioning of the search bar. In the generated version, the search bar is incorrectly placed at the bottom of the property cards. In contrast, the ground truth image clearly shows the search bar embedded within the middle property card. Although our Waffle-enhanced model correctly identified that the search bar should be horizontally centered within the UI, it failed to accurately replicate its embedded placement within the designated property card.
This specific layout mistake could potentially be rectified by applying CSS rules such as position:absolute; top:50%; left:50%; transform:translate(-50%, -50%) to the search button’s styling. The occurrence of such an error suggests that incorporating a refinement step for the generated results could be a promising avenue for future research, potentially allowing for corrections of these types of spatial misplacements.
The analysis of these failure cases offers insights into the current limitations of Waffle, particularly in areas such as precise text recognition, complex styling fidelity, and exact component placement. These observations highlight potential directions for future improvements.
A.8 Infrastructure
@@
A.9 Potential Risk and Impact
This work aims to develop a multi-modality model to help front-end developers write and understand HTML and CSS code. We assume no risk of this approach being misused.